### 09 Jan 2024 v3

@@ Abstract
Forman-Ricci curvature (FRC) has become a potent and powerful for analysing empirical networks, as the distribution of the curvature values can identify some structural information that is not readily detected by other schemes. For a graph, it is easy to compute, but it is not sensitive to higher-order structural information. That issue can be handled by going to the clique complex of the graph, and in fact, such complexes also emerge as Vietoris-Rips complexes in Topological Data Analysis. But then, with standard methods, it becomes much more computationally expensive. In this paper, we therefore develop a novel set-theoretic formulation for computing such high-order FRC in complex networks. We provide a pseudo-code, a software implementation coined FastForman, as well as a benchmark comparison with alternative implementations. Crucially, our representation theory reveals previous computational bottlenecks and also accelerates the computation of FRC.

### 24 Jan 2024 v4

@@ Abstract
Forman-Ricci curvature (FRC) is a potent and powerful tool for analysing empirical networks, as the distribution of the curvature values can identify structural information that is not readily detected by other geometrical methods. Crucially, FRC captures higher-order structural information of clique complexes of a graph or Vietoris-Rips complexes, which is not readily accessible to alternative methods. However, existing FRC platforms are prohibitively computationally expensive. Therefore, herein we develop an efficient set-theoretic formulation for computing such high-order FRC in complex networks. Significantly, our set theory representation reveals previous computational bottlenecks and also accelerates the computation of FRC. Finally, We provide a pseudo-code, a software implementation coined FastForman, as well as a benchmark comparison with alternative implementations. We envisage that FastForman will be used in Topological and Geometrical Data analysis for high-dimensional complex data sets.
@@ 7 Demonstrations
Proof 7.23.
From equation (18), we know that $F_{d}(\alpha)=\frac{|T_{\alpha}|}{(d+1)}+(d+1)-|P_{\alpha}|$. Also, recall that the FRC reaches its minimum when all neighbours of $\alpha$ are parallel, i.e., when $P_{\alpha}=N_{\alpha}$, and then, $F_{d}(\alpha)=(d+1)-|N_{\alpha}|$. On the other hand, the FRC reaches its maximum when all neighbours of $\alpha$ are transverse, $T_{\alpha}=N_{\alpha}$, where in this case $|N_{\alpha}|$ is a multiple of $(d+1)$, leading us to $F_{d}(\alpha)=\frac{|N_{\alpha}|}{(d+1)}+(d+1)$. Otherwise, $|T_{\alpha}|<|N_{\alpha}|$ and $F_{d}(\alpha)\leq\left\lfloor\frac{|N_{\alpha}|}{(d+1)}\right\rfloor+(d+1)$. In the special case where $|N_{\alpha}|=|V|-(d+1)$ and $N_{\alpha}=T_{\alpha}$ (or $N_{\alpha}=P_{\alpha}$), the upper (or lower) bound is the same as in Theorem 7.21.